OpenStudy (anonymous):

6. Solve the proportion 2/3=8/P A. P=14 B p=12 C. P=48 D. P=50 @zubhanwc3

OpenStudy (studygurl14):

2/3 = 8/OP 3(2/3) = (8/P)3 2 = 24/P P(2) = (24/P)P 2P = 24 Can you do the rest?

OpenStudy (anonymous):

P=12

OpenStudy (topdog):

12 bae

OpenStudy (anonymous):

um??

OpenStudy (studygurl14):

um what @keke_luvs_u

OpenStudy (anonymous):

u got it

OpenStudy (anonymous):

lol nun i said that to topdog when he/she said bae

OpenStudy (anonymous):

oh okay well the answer is p=12

OpenStudy (anonymous):

CAN YOU GUYS HELP WITH ONE MORE?

OpenStudy (anonymous):

what is the unit price for one CD if CDs cost 30.00? A. $9.00 B. $8.00 C. $10.00 D. $9.50

OpenStudy (anonymous):

i think is will be $10.00 cause if 3 cds cost 30.00 and 2 cds cost 20.00 so 1 will cost 10.00

OpenStudy (studygurl14):

How many CDs cost 30? It doesn't say...